In attempt to increase revenue and profits, a firm is considering a 4 percent increase in price and 11 percent increase in advertising. If the price elasticity of demand is -1.5 and the advertising elasticity of demand is +0.6, would you expect an increase or decrease in total revenues?


1
Expert's answer
2021-11-17T10:01:03-0500

Price elasticity of demand (PED) = -1.5

% increase in price = 4%


"PED=\frac{\% change indemand}{\%change in price}"


"-1.5=\frac{\% change indemand}{0.04}"


% change in demand = -1.5*0.04 = -6%

This means 4% increase in price will result as 6% decrease in demand.


Now, estimate impact of increase in advertising expenditure.

Advertising elasticity of demand (AED) = 0.6

% increase in advertising expense = 11%


"AED=\frac{\% change indemand}{\%change in advertising expense}"


"0.6=\frac{\% change indemand}{0.11}"


% change in demand = 0.6*0.11 = 6.6%

This means 11% increase in advertising expenditure will result as 6.6% decrease in demand.

Based on this, overall it can say that total revenue will increase by 0.6% in case of firm is conducting a 4 percent increase in price and 11 percent increase in advertising.